OpenStudy (anonymous):

PLEASE HELP Find the absolute minimum value of f on the given interval. f(x) = ln(x^2 + 3x + 4), [−2, 2]

OpenStudy (anonymous):

find derivative of f(x) then put f'(x) = 0

OpenStudy (anonymous):

@Nazym are u here?

OpenStudy (anonymous):

I got a wrong anser. -3/2

OpenStudy (anonymous):

it is not a simple one calculate f'(x)

OpenStudy (anonymous):

\[f^'(x)=\frac{ 2x+3 }{ x^2+3x+4 }\]

OpenStudy (anonymous):

\[x=\frac{ -3 }{ 2 }\]

OpenStudy (anonymous):

@Jonask she tell before that -3/2 is wrong answer

OpenStudy (anonymous):

I got the same answer...Didn't work

OpenStudy (anonymous):

calculate f''(x) and put x=-3/2 check the value whether -ve or +ve

OpenStudy (anonymous):

\[\left| -3/2 \right|\]

OpenStudy (anonymous):

3/2=1,5

OpenStudy (anonymous):

Jonask, it is the same thing...That's wrong(

OpenStudy (anonymous):

try to calculate f''(x)

OpenStudy (anonymous):

\[f(3/2)\]

OpenStudy (anonymous):